You are here: Foswiki>Commissioning Web>Logs>24Mar013_Log (25 Mar 2013, JuanGuerra)Edit Attach

UT 20130323 AO#2, AO#1

Observer: D.Miller (LBTO),J. Christou(LBTO),JC,Guerra(LBTO)
Telescope Operator: Geno.B.(LBTO)
SW Support: (Tucson)
AO Support: (LBTO)
Telescope Support: Jeff U (LBTO)
Instrument Support: Elliot S (LBTO)

Summary.

  • AO#2_SX and AO#_DX have been tested to move the stages and I couldn't be moved.
  • Simone sent an email to do some checks.
  • Active optics collecting data.
  • seeing good variable; 0.8" to 1.2"
  • Moon: 96% full.
  • Simone is investigating with the Arcetri team the possible causes for the stage failed. More detail in Julian's report.

Details.

The both bayside stage are not moving in the Y direction, it has been tried to push them when a command is applied, but it is showing a non response. the positions where the stages remains is showed in the photos attached. Both units are parked in the same position.

Due to the problem with the Adaptive Optics systems, the Active Optics is taking the night to collect data.

-- JuanGuerra - 25 Mar 2013
Topic revision: r1 - 25 Mar 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback